# Distributed AI Agents for Cognitive Underwater Robot Autonomy

### v1

@@ I Introduction
Within this evolving landscape, the rapid advancement of Large Language Models (LLMs) and Vision Language Models (VLMs) offers a compelling pathway toward more flexible, adaptive, and robust robot autonomy. Unlike conventional software, which executes a rigid set of programmed instructions, AI agents are designed to understand complex information, reason about it, and generate a sequence of actions to perform specific tasks [5, 6]. This approach fundamentally shifts the development paradigm from imperative programming, where every step of a task is explicitly coded (e.g., move forward 1 meter, turn left 90 degrees), to declarative goal setting, allowing the agent to autonomously determine how to fulfill its mission (e.g., inspect the valve). By leveraging vast pre-trained knowledge repositories, AI agents exhibit emergent reasoning capabilities that can adapt to new tasks and conditions without requiring exhaustive manual reprogramming [7, 8]. However, the integration of these agents in robotic applications introduces new challenges, such as ensuring robustness and verifiability of the decision made and protecting against possible hallucinations [9, 10]. The increasing capabilities of advanced AI systems also bring to the forefront critical considerations related to AI safety and verification, encompassing complex issues such as alignment of the mental models of the operators and the AI system [11, 12, 13].
UROSA is inspired by ROSA [11], which is used as a parser from high level natural language to generate actions through ROS messages. Unlike ROSA, our work integrates agentic AI at multiple levels of the autonomy architecture, able to make decisions as well as communicate through ROS. This is a novel and highly capable autonomy framework that fundamentally replaces the traditional paradigm of a human-governed main program with a distributed network of specialised and dynamically adaptable AI agents. This distributed architecture represents a radical shift towards true autonomy, where once a mission is initiated with high-level commands, the system operates with minimal human interaction and oversight. The distributed AI agents communicate and solve problems together, without needing human intervention or the development of coding or extra programmatic loops. This allows for emergent system-level intelligence, where human interaction primarily involves providing high-level commands through natural language. Each agent is responsible for a specific aspect of the robot’s operational workflow, ranging from multimodal perception (e.g., vision, depth, sonar) to high-level strategic planning. Crucially, the AI agents within UROSA are not merely generative; they are designed as agentic AI entities, capable of autonomous decision-making and action-taking within their environment without requiring continuous human intervention in the loop [14, 15].
@@ V-A3 Flexible Motion Planning for a UVMS
To provide a quantitative benchmark, we compared its performance against a baseline traditional planner using a naive A* search on a Probabilistic Roadmap (PRM) [60]. To test the reliability of the AI’s language understanding, we conducted 10 trials using prompts with varied phrasing and intent. These included explicit instructions for avoidance (e.g., Go to goal 1 and check the obstacles on the way) as well as more ambiguous phrasing (e.g., Inspect goal 2 and check for any boxes on your way). Crucially, the test also included direct commands that omitted any mention of obstacle avoidance, such as Inspect goal 2. In these cases, to test literal command adherence, the Commander AI Agent intentionally disregarded perceptual data about obstacles and commanded a direct path, leading to a planned collision. The averaged results of these trials are summarized in Table III.
@@ V-B2 Online Behavioural Tuning via Teacher-Student Interaction
To demonstrate more direct online behavioural adaptation, we validated our novel Teacher-Student Instructional Tuning mechanism. In this paradigm, a Teacher agent guides a Student agent’s policy by generating a new, more restrictive SYSTEM prompt as a form of instructive feedback, creating a powerful meta-learning loop. We conducted a series of real-world experiments in a cluttered tank environment (Fig. 4) where the Commander AI Agent acted as the Teacher and the Perception & Scene Reasoning AI Agent was the Student.
@@ V-C Dynamic, Predictive System Diagnostics
This capability was evaluated through a series of proof-of-concept experiments where we simulated various hardware failure modes. For each test case, we programmatically disabled specific thrusters via the vehicle’s control interface. The ArduPilot system [61], observing the lack of response, would then generate an updated vehicle status JSON reflecting the fault. The Diagnostic AI Agent’s task was to monitor this stream and correctly identify which thrusters were malfunctioning. We ran 5 trials for each fault configuration, and the agent correctly diagnosed the system state with 100% accuracy across all tests, as summarised in Table VII. To further demonstrate the agent’s ability to track a dynamically changing system state, we conducted a specific transition test. First, we disabled thrusters 2 and 3 to induce a fault. After the agent correctly diagnosed this failure, we re-enabled the thrusters in the interface. The agent’s subsequent report immediately and accurately reflected the system’s return to a healthy state. Figure 8 shows the agent’s real-time terminal output, capturing these precise system faults.

### v2

@@ I Introduction
Within this evolving landscape, the rapid advancement of Large Language Models (LLMs) and Vision Language Models (VLMs) offers a compelling pathway toward more flexible, adaptive, and robust robot autonomy. Unlike conventional software, which executes a rigid set of programmed instructions, AI agents are designed to understand complex information, reason about it, and generate a sequence of actions to perform specific tasks [5, 6]. This approach fundamentally shifts the development paradigm from imperative programming, where every step of a task is explicitly coded (e.g., move forward 1 meter, turn left 90 degrees), to declarative goal setting, allowing the agent to autonomously determine how to fulfill its mission (e.g., inspect the valve). By leveraging vast pre-trained knowledge repositories, AI agents exhibit emergent reasoning capabilities that can adapt to new tasks and conditions without requiring exhaustive manual reprogramming [7, 8]. However, the integration of these agents in robotic applications introduces new challenges, such as ensuring robustness and verifiability of the decision made and protecting against possible hallucinations [9, 10]. The increasing capabilities of advanced AI systems also bring to the forefront critical considerations related to AI safety and verification, encompassing complex issues such as alignment of the mental models of the operators and the AI system [11].
UROSA is inspired by ROSA [12], which is used as a parser from high level natural language to generate actions through ROS messages. Unlike ROSA, our work integrates agentic AI at multiple levels of the autonomy architecture, able to make decisions as well as communicate through ROS. This is a novel and highly capable autonomy framework that fundamentally replaces the traditional paradigm of a human-governed main program with a distributed network of specialised and dynamically adaptable AI agents. This distributed architecture represents a radical shift towards true autonomy, where once a mission is initiated with high-level commands, the system operates with minimal human interaction and oversight. The distributed AI agents communicate and solve problems together, without needing human intervention or the development of coding or extra programmatic loops. This allows for emergent system-level intelligence, where human interaction primarily involves providing high-level commands through natural language. Each agent is responsible for a specific aspect of the robot’s operational workflow, ranging from multimodal perception (e.g., vision, depth, sonar) to high-level strategic planning. Crucially, the AI agents within UROSA are not merely generative; they are designed as agentic AI entities, capable of autonomous decision-making and action-taking within their environment without requiring continuous human intervention in the loop [13].
@@ V-A3 Flexible Motion Planning for a UVMS
To provide a quantitative benchmark, we compared its performance against a baseline traditional planner using a naive A* search on a Probabilistic Roadmap (PRM). To test the reliability of the AI’s language understanding, we conducted 10 trials using prompts with varied phrasing and intent. These included explicit instructions for avoidance (e.g., Go to goal 1 and check the obstacles on the way) as well as more ambiguous phrasing (e.g., Inspect goal 2 and check for any boxes on your way). Crucially, the test also included direct commands that omitted any mention of obstacle avoidance, such as Inspect goal 2. In these cases, to test literal command adherence, the Commander AI Agent intentionally disregarded perceptual data about obstacles and commanded a direct path, leading to a planned collision. The averaged results of these trials are summarized in Table III.
@@ V-B2 Online Behavioural Tuning via Teacher-Student Interaction
To demonstrate more direct online behavioural adaptation, we validated our novel Teacher-Student Instructional Tuning mechanism. In this paradigm, a Teacher agent guides a Student agent’s policy by generating a new, more restrictive SYSTEM prompt as a form of instructive feedback, creating a powerful meta-learning loop. We conducted a series of real-world experiments in a cluttered tank environment, shown in Fig. 4, where the Commander AI Agent acted as the Teacher and the Perception & Scene Reasoning AI Agent was the Student.
@@ V-C Dynamic, Predictive System Diagnostics
This capability was evaluated through a series of proof-of-concept experiments where we simulated various hardware failure modes. For each test case, we programmatically disabled specific thrusters via the vehicle’s control interface. The ArduPilot system [58] communicating with Stonefish [59, 60], observing the lack of response, would then generate an updated vehicle status JSON reflecting the fault. The Diagnostic AI Agent’s task was to monitor this stream and correctly identify which thrusters were malfunctioning. We ran 5 trials for each fault configuration, and the agent correctly diagnosed the system state with 100% accuracy across all tests, as summarised in Table VII. To further demonstrate the agent’s ability to track a dynamically changing system state, we conducted a specific transition test. First, we disabled thrusters 2 and 3 to induce a fault. After the agent correctly diagnosed this failure, we re-enabled the thrusters in the interface. The agent’s subsequent report immediately and accurately reflected the system’s return to a healthy state. Figure 8 shows the agent’s real-time terminal output, capturing these precise system faults.